Determine the polynomial function whose graph passes through the points (2, 4), (3,6) and (5,10). Also sketch the graph of the polynomial function. (Using Cramers Method).
1
Expert's answer
2020-12-15T19:42:33-0500

we have three points so we use second polynomial fitting like that

"p(x) = a + bx + c^2"

the points are (2, 4); (3, 6) and (5, 10).

"y_1 =" "ax_1^2 + bx_1+c"

"y_2 = ax_2^2 + bx_2 + c"

"y_3 = ax_3^2 + bx_3 + c"

The task is to find a, b and c. Start by substituting each of the points into the equation, we have

4 = a(2)2 + b(2) + c;

6 = a(3)2 + b(3) + c;

10 = a(5)2 + b(5) + c;

We can write this more compactly as a matrix equation

"\begin{bmatrix}
 4&2&1\\
 9&3&1\\
25&5&1
\end{bmatrix}" "\begin{bmatrix}
 a\\
b\\
c

\end{bmatrix}" = "\begin{bmatrix}
 4\\
 6\\
10
\end{bmatrix}"


Or we use Cramer's rule for:

4a + 2b + c = 4

9a + 3b + c = 6

25a + 5b + c = 10


D = "\begin{bmatrix}
 4&2&1\\
 9&3&1\\
25&5&1
\end{bmatrix}"= (4*3*1 + 2*1*25 + 1*9*5) - (25*3*1 + 5*1*4 + 1*9*2) = -6;


Dx = "\begin{bmatrix}
 4&2&1\\
 6&3&1\\
10&5&1
\end{bmatrix}" = (4*3*1 + 2*1*10 + 1*6*5) - (10 * 3 *1 + 5*1*4 + 1*6*2) = 0;


x = "\frac{D_x}{D} = \frac{0}{-6} = 0"


Dy = "\begin{bmatrix}
 4&4&1\\
 9&6&1\\
25&10&1
\end{bmatrix}" = (4*6*1 + 4*1*25 + 1*9*10) - (25*6 + 10*1*4+1*9*4) = -12


"b = \frac{D_y}{D} = \frac{-12}{--6} = 2"


D"_z = \begin{bmatrix}
 4&2&4\\
 9&3&6\\
25&5&10\end{bmatrix}" = 0

"c = \frac{D_z}{D} = \frac{0}{-12} = 0"

4 + c = 4 -> c = 0;

y = 2x